OpenStudy (teller):

Someone mind helping me ? Consider the line y=-3/5x+5 Find the equation of the line that is parallel to this line and passes through the point (−9, -6) Find the equation of the line that is perpendicular to this line and passes through the point (−9,-6).

OpenStudy (math&ing001):

We know parallel lines have the same slope, using the formula y-y1=m(x-x1) with (x1,y1) being the point given you can find the equation of your line.

OpenStudy (teller):

Huh?

OpenStudy (teller):

I'm a little confused...

OpenStudy (math&ing001):

OK, let's start with the A, B, C

OpenStudy (math&ing001):

Do you know what a slope is ?

OpenStudy (teller):

Not really..

OpenStudy (math&ing001):

OK, here's an equation of a line in standard form y = m * x + b m is the slope and b is intercept.

OpenStudy (math&ing001):

Can you identify the slope in your equation ?

OpenStudy (teller):

is it -3/5?

OpenStudy (math&ing001):

Yep !

OpenStudy (teller):

so it's y=-3/5x times 5?

OpenStudy (math&ing001):

No that's it, slope here is m=-3/5 you got it. Now we need to find the equation of the line parallel to y=-3/5x+5 that passes through (-9,-6)

OpenStudy (teller):

So how do we do that?

OpenStudy (math&ing001):

We know parallel lines have the same slope, so the the line we are looking for will also have a slope of -3/5, plus we know that it passes through the point (-9,-6).

OpenStudy (math&ing001):

There's a formula that helps find a line's equation using (slope+point) which is : y-y1=m(x-x1) y and x are the variables of the equation m is the slope (x1,y1) are the coordinates of the point which the line crosses

OpenStudy (teller):

So what do I do ?

OpenStudy (math&ing001):

Plug in the values in the equation above.

OpenStudy (teller):

I'm confused tho I know where -3/5 goes but where do (-9,-6) go?

OpenStudy (math&ing001):

Hint : (x1,y1) are the coordinates of the point which the line crosses

OpenStudy (math&ing001):

which the line passes through*

OpenStudy (teller):

So it would be -9-(-6)=-3/5(x-x1)?

OpenStudy (math&ing001):

Not quite, x and y are variables of the equation. You only plug in x1 and y1.

OpenStudy (teller):

so y-(-9)=-3/5(x-(-6)??

OpenStudy (math&ing001):

x1 is the x-coordinate of the point y1 is the y-coordinate of the point

OpenStudy (teller):

I'm confused.

OpenStudy (math&ing001):

The coordinates are always written as (x-coordinate, y-coordinate) in the Cartesian system.

OpenStudy (teller):

I'm sorry but I have to go .Thanks for the help anyways.

OpenStudy (math&ing001):